# Structure-Enhanced Protein Instruction Tuning: Towards General-Purpose Protein Understanding

### v2

@@ 4 Structure-Enhanced Protein Instruction Tuning
4.1 Model Architecture
In this subsection, we introduce three main components of the SEPIT framework: 1) a sequence/structure fused protein encoder, 2) a linear projector, and 3) a large language model with mixture of experts modules. The illustration of the model architecture is depicted in Figure 1.
Figure 1: (a) The model architecture of the SEPIT framework includes sequence/structure fused protein encoder, linear projector, and LLMs with MoEs modules, (b) example of instruction format.
@@ 4.1 Model Architecture
Considering the abundance of sequence-only data and the relatively small amount of sequence-structure paired data (whether experimentally-determined structures or computed structures) within the protein domain (Varadi et al., 2022; Steinegger et al., 2019; Steinegger & Söding, 2018; Suzek et al., 2015), we propose a Sequence/Structure Fused Protein Encoder that is capable of accommodating inputs in either form. Additionally, through leveraging the limited sequence-structure paired data, we aim to enhance the model’s performance when dealing with sequence-only inputs.
For sequence-only data, numerous pLMs (Rives et al., 2021; Lin et al., 2023; Brandes et al., 2022; Elnaggar et al., 2022; Rao et al., 2021) have already been pre-trained on it. To enable them to perceive structural information, we have designed structure-aware modules for pLMs. Mainstream pLMs, such as ESM (Rives et al., 2021; Lin et al., 2023), are encoder-only architectures, consisting of multiple Transformer encoder layers, which primarily comprise self-attention modules (Vaswani et al., 2017) and feed-forward network (FFN). Here, our main focus is on the self-attention modules (for simplicity, we discuss the scenario with single-head and assume that the dimensions of the query, key, and value are all equal to the hidden size $d$). Let $\boldsymbol{X}^{(l)}=[\boldsymbol{x}_{1}^{(l)},\boldsymbol{x}_{2}^{(l)},\cdots,\boldsymbol{x}_{N}^{(l)}]^{\top}$ denote the input to self-attention module in $l$-th layer, where $\boldsymbol{x}_{i}^{(l)}\in\mathbb{R}^{d}$ is the $d$-dimension representation of the $i$-th residue out of the $N$ residues in the protein. The self-attention module then works as follows:
$\boldsymbol{A}^{(l)}=$ $\frac{\boldsymbol{X}^{(l)}\boldsymbol{W}_{Q}^{(l)}(\boldsymbol{X}^{(l)}\boldsymbol{W}^{(l)}_{K})^{\top}}{\sqrt{d}},$ (1)
$\operatorname{Attn}(\boldsymbol{X}^{(l)})=$ $\operatorname{softmax}(\boldsymbol{A}^{(l)})\boldsymbol{X}^{(l)}\boldsymbol{W}^{(l)}_{V},$ (2)
@@
$\boldsymbol{\Delta}=\phi\left(\boldsymbol{\psi}_{(i,j)}\boldsymbol{W}_{a}\right)\boldsymbol{W}_{b},$ (3)
@@
$\psi_{(i,j)}^{k}=-\frac{1}{\sqrt{2\pi}|\sigma^{k}|}\exp\left(-\frac{1}{2}\left(\frac{\|\boldsymbol{c}_{i}-\boldsymbol{c}_{j}\|-\mu^{k}}{|\sigma^{k}|}\right)^{2}\right),\;k=1,...,K,$ (4)
@@
$\hat{\boldsymbol{A}}^{(l)}=\boldsymbol{A}^{(l)}+\boldsymbol{\Delta},$ (5)
and added as structure positional encoding to the embedding input of pLMs:
$\hat{\boldsymbol{X}}^{(0)}=\boldsymbol{X}^{(0)}+\omega\left(\sum_{j\in[n]}\boldsymbol{\psi}_{(i,j)}\right)\boldsymbol{W}_{c},$ (6)
where $\omega$ signifies the coefficient that regulates the magnitude of the structure positional encoding and $\boldsymbol{W}_{c}\in\mathbb{R}^{K\times d}$ is learnable linear transformation. It is noteworthy that when the input to the sequence/structure fused protein encoder consists solely of the protein sequence, lacking structural information, the structure-aware module will be automatically disabled. This allows it to adapt to different types of protein inputs.
@@
To bridge proteins with natural language, a module is required to link the protein encoder and the LLMs decoder. Prior work in the MLLMs field has contributed outstanding methods such as Q-former (Li et al., 2023), linear projector (Liu et al., 2024a), and merging tokens before the linear projector (Zhu et al., 2023). Considering the vast differences between proteins and visual images - that is, the former requires the retention of more information of all residues (as any change in the amino acid sequence can lead to significant structural differences, resulting in profoundly different properties and functions), whereas the latter possesses some degree of information redundancy - we opt for a simple linear projector to reduce information loss.
@@
Due to the understanding of proteins being a complex multi-task problem, the various properties and functions of proteins can exhibit significant changes with subtle variations in the amino acid sequence. As the hub of "understanding" within the entire framework, the capabilities of LLMs are crucial. Previous scaling laws (Kaplan et al., 2020) have suggested that larger parameter sizes can endow LLMs with stronger capabilities; however, the additional computation costs brought about by increased activated parameters are intolerable for us. Therefore, we seek to leverage mixture-of-experts (MoEs) to achieve higher parameter sizes without increasing the number of activated parameters, thereby enhancing the model’s capacity and generalization ability. In our framework, the MoEs module replaces the FFN module in each Transformer decoder layer. The MoEs module works as follows (Lepikhin et al., 2020; Jacobs et al., 1991; Zoph et al., 2022; Lin et al., 2024):
$y=\sum_{i=1}^{n}\text{G}(\boldsymbol{x})_{i}\cdot\text{E}_{i}(\boldsymbol{x}),$ (7)
Here, $\boldsymbol{x}$ is assumed to be the input to the original FFN layer, and E represents the $n$ experts in the MoEs, each of which has the exact same structure as the original FFN layer. G represents the gating network, $\text{G}(\boldsymbol{x})_{i}$ denotes the gate weight for the $i$-th expert, and $\text{E}_{i}(\boldsymbol{x})$ is the output of the $i$-th expert. For the gating network, we employed the commonly used linear TopK gate:
$\text{G}(\boldsymbol{x}):=\operatorname{Softmax}\left(\operatorname{TopK}\left(\boldsymbol{x}\cdot\boldsymbol{W}_{\text{G}}\right)\right),$ (8)
and we imposed auxiliary loss to ensure the token balance among the experts (Zoph et al., 2022; Lepikhin et al., 2020):
$\mathcal{L}_{\text{aux}}=n\cdot\sum_{i=1}^{n}f_{i}\cdot P_{i},$ (9)
@@
4.2 Training Pipeline
Figure 2: The three-stage training pipeline of SEPIT with a warm-up stage (Stage 0) for protein encoder, and a two-stage instruction tuning (Stage 1 & Stage 2).
@@ 4.2 Training Pipeline
In this stage, our primary goal is to warm up our protein encoder. Although the pLM is already pre-trained, the structure-aware module we plug in is randomly initialized. To address this, we leverage two main pre-training paradigms. Firstly, to enable the structure-aware module to better perceive structural information, we follow the common practice in molecular self-supervised learning (Godwin et al., 2022; Zaidi et al., 2023; Luo et al., 2023a; Zhou et al., 2023), which involves structure denoising tasks. For the input 3D coordinates of protein residues $\boldsymbol{C}=[\boldsymbol{c}_{1},\boldsymbol{c}_{2},\cdots,\boldsymbol{c}_{N}]^{\top}$, we apply noise to obtain nosied coordinates $\tilde{\boldsymbol{C}}=[\tilde{\boldsymbol{c}}_{1},\tilde{\boldsymbol{c}}_{2},\cdots,\tilde{\boldsymbol{c}}_{N}]^{\top}$, where $\tilde{\boldsymbol{c}}_{i}=\boldsymbol{c}_{i}+\alpha\boldsymbol{\delta}_{i},\;\boldsymbol{\delta}_{i}\sim\mathcal{N}(0,I)$, $\alpha$ is a scaler used to control the magnitude of noise. Then we predict the applied noise based on them (equivalent to predicting the original 3D coordinates). The formula for denoise loss is as follows:
$\mathcal{L}_{\text{Denoise}}=\frac{1}{3n}\sum_{i=1}^{n}\sum_{j=1}^{3}\left(\boldsymbol{\delta}_{i}^{j}-\hat{\boldsymbol{\delta}_{i}^{j}}\right)^{2},$ (10)
@@
$\hat{\boldsymbol{\delta}}_{i}^{j}=\left(\sum_{i=1}^{n}\boldsymbol{\Delta}_{ik}\boldsymbol{D}_{ik}^{j}\boldsymbol{X}^{(L+1)}_{k}\boldsymbol{W}_{m}\right)\boldsymbol{W}_{n},\quad\boldsymbol{D}_{ik}=\frac{\mathbf{c}_{i}-\mathbf{c}_{k}}{\|\mathbf{c}_{i}-\mathbf{c}_{k}\|}.$ (11)
@@
$\mathcal{L}_{\text{CLIP}}=-\frac{1}{2B}\sum_{i=1}^{B}\left(\log\frac{\exp(\boldsymbol{p}_{i}\cdot\boldsymbol{t}_{i}/\tau)}{\sum_{j=1}^{B}\exp(\boldsymbol{p}_{i}\cdot\boldsymbol{t}_{j}/\tau)}+\log\frac{\exp(\boldsymbol{p}_{i}\cdot\boldsymbol{t}_{i}/\tau)}{\sum_{j=1}^{B}\exp(\boldsymbol{p}_{j}\cdot\boldsymbol{t}_{i}/\tau)}\right),$ (12)
@@
$\mathcal{L}_{\text{Stage 0}}=\mathcal{L}_{\text{Denoise}}+\mathcal{L}_{\text{CLIP}}+\mathcal{L}_{\text{MLM}}.$ (13)
@@
In this stage, our fundamental objective is to further align proteins with their natural language descriptions through the paradigm of conditional generation (Liu et al., 2024a; Ouyang et al., 2022), utilizing protein caption instructions. To ensure consistency in the model’s handling of different forms of protein inputs, we randomly input protein data, both those with only sequences and those paired with structures, into our protein encoder at probabilities of $15\%$ and $85\%$, respectively. The output protein representation sequences are mapped to the textual space through linear projector, in conjunction with protein caption instructions to guide LLMs in producing straightforward descriptions of proteins, such as function, family, subcellular localization, and overall descriptions. Formally, consider a protein-text pair $(\mathcal{P},\mathcal{T})$ similar to that in Stage 0, given the output sequence of the protein encoder $\boldsymbol{S_{p}}$, and instructions $\boldsymbol{S_{\text{instruct}}}$, the objective of Stage 1 is as follows:
$\mathcal{L}_{\text{Stage 1}}=p(\boldsymbol{S_{t}}|f(\mathbf{S_{p}}),\boldsymbol{S_{\text{instruct}}})=\prod_{i=1}^{L}p_{\boldsymbol{\theta}}(s_{i}|\boldsymbol{S_{p}},\boldsymbol{S}_{\text{instruct},<i},\boldsymbol{S}_{\boldsymbol{t},<i}),$ (14)
@@
$\mathcal{L}_{\text{Stage 2}}=\mathcal{L}_{\text{Stage 1}}+\beta\mathcal{L}_{\text{aux}},$ (15)
@@
5 Experiments
In this section, we will first introduce the experimental setting in this paper. Subsequently, we will comprehensively demonstrate the effectiveness of SEPIT and its various designs through performance comparisons and ablation studies. Finally, we will delve deeper into the characteristics of SEPIT through case studies.
5.1 Experimental Setting
First, we provide a brief overview of the main settings in experiments. More detailed information such as implementation details and dataset construction details can be found in Appendix B.1 and A.

### v3

@@ 4. Structure-Enhanced Instruction Tuning
4.1. Model Architecture
In this subsection, we introduce three main components of the SEPIT framework: 1) a sequence/structure fused protein encoder, 2) a linear projector, and 3) a large language model with MoEs modules. The model architecture is shown in Figure 1.
@@ 4.1. Model Architecture
Given the vast availability of sequence-only data and limited sequence-structure paired data (from both experimental and computational sources) within protein domain (Varadi et al., 2022; Steinegger et al., 2019; Steinegger and Söding, 2018; Suzek et al., 2015), we propose a sequence/ structure fused protein encoder that is capable of accommodating inputs in either form. Additionally, through leveraging the scarce sequence-structure paired data, we aim to enhance the model’s performance when dealing with sequence-only inputs.
For sequence-only data, numerous pLMs (Rives et al., 2021; Lin et al., 2023; Brandes et al., 2022; Elnaggar et al., 2022; Rao et al., 2021) have already been pre-trained on it. To facilitate their perception of structural information, we have designed a structure-aware module for pLMs. This module encodes the 3D structural information into the attention matrix and positional encoding of pLMs in a simple but effective way. Specifically, the attention mechanism of pLMs is defined as follows (for simplicity, we discuss the scenario with single-head and assume that the dimensions of the query, key, and value are all equal to the hidden size $d$). Let $\boldsymbol{X}^{(l)}=[\boldsymbol{x}_{1}^{(l)},\boldsymbol{x}_{2}^{(l)},\cdots,\boldsymbol{x}_{N}^{(l)}]^{\top}$ denote the input to self-attention module in $l$-th layer, where $\boldsymbol{x}_{i}^{(l)}\in\mathbb{R}^{d}$ is the $d$-dimension representation of the $i$-th residue out of the $N$ residues in the protein. The self-attention module then works as follows:
(1) $\boldsymbol{A}^{(l)}=$ $\frac{\boldsymbol{X}^{(l)}\boldsymbol{W}_{Q}^{(l)}(\boldsymbol{X}^{(l)}\boldsymbol{W}^{(l)}_{K})^{\top}}{\sqrt{d}},$
(2) $\operatorname{Attn}(\boldsymbol{X}^{(l)})=$ $\operatorname{softmax}(\boldsymbol{A}^{(l)})\boldsymbol{X}^{(l)}\boldsymbol{W}^{(l)}_{V},$
@@
(3) $\boldsymbol{\Delta}=\phi\left(\boldsymbol{\psi}_{(i,j)}\boldsymbol{W}_{a}\right)\boldsymbol{W}_{b},$
@@
(4) $\psi_{(i,j)}^{k}=-\frac{1}{\sqrt{2\pi}|\sigma^{k}|}\exp\left(-\frac{1}{2}\left(\frac{\|\boldsymbol{c}_{i}-\boldsymbol{c}_{j}\|-\mu^{k}}{|\sigma^{k}|}\right)^{2}\right),\;k=1,...,K,$
@@
(5) $\hat{\boldsymbol{A}}^{(l)}=\boldsymbol{A}^{(l)}+\boldsymbol{\Delta},$
and added as structure positional encoding to the input of pLMs:
(6) $\hat{\boldsymbol{X}}^{(0)}=\boldsymbol{X}^{(0)}+\omega\left(\sum_{j\in[n]}\boldsymbol{\psi}_{(i,j)}\right)\boldsymbol{W}_{c},$
where $\omega$ signifies the coefficient that regulates the magnitude of the structure positional encoding and $\boldsymbol{W}_{c}\in\mathbb{R}^{K\times d}$ is learnable linear transformation. It is noteworthy that when the input to the sequence/structure fused protein encoder consists solely of the protein sequence, lacking structural information, the structure-aware module will be automatically disabled. This allows it to adapt to different protein input formats.
Figure 2. The three-stage training pipeline of SEPIT with a warm-up stage (stage 0) for protein encoder and two instruction tuning stages (stage 1 & stage 2).
@@
To bridge proteins with natural language, a module is required to link the protein encoder and the LLMs decoder. Prior work in the MLLMs field has contributed outstanding methods such as Q-former (Li et al., 2023), linear projector (Liu et al., 2024a), and merging tokens before the linear projector (Zhu et al., 2024). Considering the vast differences between proteins and visual images - that is, the former requires the retention of more information of all residues (as any change in the amino acid sequence can lead to significant structural differences, resulting in profoundly different properties and functions), whereas the latter possesses some degree of information redundancy - we opt for a simple linear projector to reduce information loss.
@@
Due to the understanding of proteins being a complex multi-task problem, the various properties and functions of proteins can exhibit significant changes with subtle variations in the amino acid sequence. As the hub of ”understanding” within the entire framework, the capabilities of LLMs are crucial. Previous scaling laws (Kaplan et al., 2020) have suggested that larger parameter sizes can endow LLMs with stronger capabilities; however, the additional computation costs brought about by increased activated parameters are intolerable for us. Therefore, we seek to leverage mixture-of-experts (MoEs) to achieve higher parameter sizes without increasing the number of activated parameters, thereby enhancing the model’s capacity and generalization ability. In our framework, the MoEs module replaces the FFN module in each Transformer decoder layer. The MoEs module works as follows (Lepikhin et al., 2020; Jacobs et al., 1991; Zoph et al., 2022; Lin et al., 2024; Zhang et al., 2025; Xiao et al., 2024):
(7) $y=\sum_{i=1}^{n_{e}}\text{G}(\boldsymbol{x})_{i}\cdot\text{E}_{i}(\boldsymbol{x}).$
Here, $\boldsymbol{x}$ is assumed to be the input to the original FFN layer, and E represents the $n_{e}$ experts in the MoEs, each of which has the exact same structure as the original FFN layer. G represents the gating network, $\text{G}(\boldsymbol{x})_{i}$ denotes the gate weight for the $i$-th expert, and $\text{E}_{i}(\boldsymbol{x})$ is the output of the $i$-th expert. For the gating network, we employ the commonly used linear TopK gate:
(8) $\text{G}(\boldsymbol{x}):=\operatorname{Softmax}\left(\operatorname{TopK}\left(\boldsymbol{x}\cdot\boldsymbol{W}_{\text{G}}\right)\right),$
and we impose auxiliary loss to ensure the token balance among the experts (Zoph et al., 2022; Lepikhin et al., 2020):
(9) $\mathcal{L}_{\text{aux}}=n\cdot\sum_{i=1}^{n}f_{i}\cdot P_{i},$
@@
4.2. Training Pipeline
@@
In this stage, our primary goal is to warm up our protein encoder. Although the pLM is already pre-trained, the structure-aware module that we plug in is randomly initialized. To address this, we leverage two main pre-training paradigms. Firstly, to enable the structure-aware module to better perceive structural information, we follow the common practice in molecular self-supervised learning (Godwin et al., 2022; Zaidi et al., 2023; Luo et al., 2023a; Zhou et al., 2023), which involves structure denoising tasks. For the input 3D coordinates of protein residues $\boldsymbol{C}=[\boldsymbol{c}_{1},\boldsymbol{c}_{2},\cdots,\boldsymbol{c}_{N}]^{\top}$, we apply noise to obtain noised coordinates $\tilde{\boldsymbol{C}}=[\tilde{\boldsymbol{c}}_{1},\tilde{\boldsymbol{c}}_{2},\cdots,\tilde{\boldsymbol{c}}_{N}]^{\top}$, where $\tilde{\boldsymbol{c}}_{i}=\boldsymbol{c}_{i}+\alpha\boldsymbol{\delta}_{i},\;\boldsymbol{\delta}_{i}\sim\mathcal{N}(0,I)$, $\alpha$ is a scaler used to control the magnitude of noise. Then we predict the applied noise based on them (equivalent to predicting the original 3D coordinates). The formula for denoise loss is as follows:
(10) $\mathcal{L}_{\text{Denoise}}=\frac{1}{3n}\sum_{i=1}^{n}\sum_{j=1}^{3}\left(\boldsymbol{\delta}_{i}^{j}-\hat{\boldsymbol{\delta}_{i}^{j}}\right)^{2},$
@@
(11) $\hat{\boldsymbol{\delta}}_{i}^{j}=\left(\sum_{i=1}^{n}\boldsymbol{\Delta}_{ik}\boldsymbol{D}_{ik}^{j}\boldsymbol{X}^{(L+1)}_{k}\boldsymbol{W}_{m}\right)\boldsymbol{W}_{n},\quad\boldsymbol{D}_{ik}=\frac{\mathbf{c}_{i}-\mathbf{c}_{k}}{\|\mathbf{c}_{i}-\mathbf{c}_{k}\|}.$
@@
(12) $\mathcal{L}_{\text{CLIP}}=-\frac{1}{2B}\sum_{i=1}^{B}\left(\log\frac{\exp(\boldsymbol{p}_{i}\cdot\boldsymbol{t}_{i}/\tau)}{\sum_{j=1}^{B}\exp(\boldsymbol{p}_{i}\cdot\boldsymbol{t}_{j}/\tau)}+\log\frac{\exp(\boldsymbol{p}_{i}\cdot\boldsymbol{t}_{i}/\tau)}{\sum_{j=1}^{B}\exp(\boldsymbol{p}_{j}\cdot\boldsymbol{t}_{i}/\tau)}\right),$
@@
(13) $\mathcal{L}_{\text{stage 0}}=\mathcal{L}_{\text{Denoise}}+\mathcal{L}_{\text{CLIP}}+\mathcal{L}_{\text{MLM}}.$
@@
Table 1. Performance comparisons on open-ended generation and closed-set answer tasks.
Model Activated Parameters Open-Ended Closed-Set BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L METEOR BERT-P BERT-R BERT-F1 Accuracy Zero-Shot GPT-3.5-turbo N/A 3.26 0.02 12.41 3.14 11.06 10.44 85.18 85.40 85.24 56.56% Claude-3-haiku N/A 3.00 0.07 12.10 2.65 10.62 9.28 86.04 85.47 85.70 59.14% GPT-4-turbo N/A 4.21 0.08 12.78 2.93 11.57 10.41 86.91 85.56 85.71 58.58% GPT-4o-mini N/A 2.36 0.01 10.72 1.93 9.03 9.00 85.08 85.38 85.19 41.58% GPT-4o N/A 3.26 0.14 15.02 3.73 12.30 13.01 85.28 86.51 85.86 59.70% DeepSeek-V3 671B 2.75 0.02 12.15 3.37 10.54 10.25 85.32 85.63 85.44 56.49% Galactica 1.3B 0.43 0.01 3.49 0.41 2.67 2.44 85.79 82.61 84.08 39.15% BioMedGPT 7B 0.83 0.01 4.90 0.49 3.26 4.59 85.51 84.95 85.14 38.61% Mol-Instructions 7B 0.53 0.01 5.96 0.39 4.64 5.51 83.81 84.41 84.06 BioT5+ 252M 3.88 1.92 12.12 4.88 10.37 14.26 85.14 85.93 85.48 InstructProtein 1.3B 5.50 2.97 14.80 5.68 13.76 13.17 85.34 85.92 85.57 48.37% Instruction Tuning TinyLlama 1.1B 51.16 43.44 65.41 51.26 62.31 60.80 93.97 94.37 94.16 74.09% OpenLlama-v2 3B 36.19 30.65 48.33 36.52 45.53 49.01 92.92 91.87 92.35 71.77% Llama2 7B 57.02 49.47 70.80 57.24 67.78 65.96 94.95 95.17 95.05 71.68% Sequence-Only Protein Instruction Tuning PIT-TinyLlama 1.8B 57.82 50.02 71.34 58.16 68.35 66.19 95.18 95.28 95.26 76.02% PIT-TinyLlama-MoEs 1.8B 57.92 50.01 72.13 58.21 69.19 66.29 95.31 95.30 95.29 78.56% Structure-Enhanced Protein Instruction Tuning SEPIT-TinyLlama 1.8B 58.43 51.04 72.34 58.77 69.13 67.91 95.32 95.59 95.44 79.05% SEPIT-Llama2 8B 60.81 52.37 74.80 60.84 71.62 68.43 95.81 95.73 95.76 79.97% SEPIT-TinyLlama-MoEs 1.8B 60.28 52.16 74.22 60.29 71.13 68.27 95.62 95.69 95.64 79.73%
@@
In this stage, our fundamental objective is to further align proteins with their natural language descriptions through the paradigm of conditional generation (Liu et al., 2024a; Ouyang et al., 2022), utilizing protein caption instructions. To ensure consistency in the model’s handling of different forms of protein inputs, we randomly input protein data, both those with only sequences and those paired with structures, into our protein encoder at probabilities of $15\%$ and $85\%$, respectively. The output protein representation sequences are mapped into the textual space through linear projector, in conjunction with protein caption instructions to guide LLMs in producing straightforward descriptions of proteins, such as function, family, subcellular localization, and overall descriptions. Formally, consider a protein-text pair $(\mathcal{P},\mathcal{T})$ similar to that in stage 0, given the output sequence of the protein encoder $\boldsymbol{S_{p}}$, and instructions $\boldsymbol{S_{\text{instruct}}}$, the objective of stage 1 is as follows:
(14) $\mathcal{L}_{\text{stage 1}}=p(\boldsymbol{S_{t}}|f(\mathbf{S_{p}}),\boldsymbol{S_{\text{instruct}}})=\prod_{i=1}^{L}p_{\boldsymbol{\theta}}(s_{i}|\boldsymbol{S_{p}},\boldsymbol{S}_{\text{instruct},<i},\boldsymbol{S}_{\boldsymbol{t},<i}),$
@@
(15) $\mathcal{L}_{\text{stage 2}}=\mathcal{L}_{\text{stage 1}}+\beta\mathcal{L}_{\text{aux}},$
@@
Table 2. Performance comparisons on overlapping properties and functions.
5. Experiments
In this section, we will first outline the experimental setting. Subsequently, we will demonstrate SEPIT’s effectiveness and design benefits via performance comparisons and ablation studies. Finally, we will delve deeper into the characteristics of SEPIT through generalization analysis and case studies. Additional analysis, including error pattern analysis are provided in the Appendix C.5.
5.1. Experimental Setting
First, we provide a brief overview of the main experimental settings, with more detailed information available in Appendix B, C.1.